One of the racks in our datacenter is currently experiencing an electrical incident affecting some of the servers in that rack. Our technical teams are investigating and working to bring the affected servers back online as soon as possible. We recommend that you refrain from stopping/rebooting your VPS servers, and these operations will be automatic when the affected VPS servers can be migrated.
Updated: 13:00 CET (12:00 GMT): The power has been restored to the affected servers. Our technical teams are investigating the cause of the outage.
A fibre cut in New Jersey is affecting one of our international circuits between Paris and Baltimore. The cut occured at just after midnight CEST (22:00 UTC). Our suppliers are working to repair the break, but the local power company must repair damaged electrical infrastructure before the fibre resplicing can commence.
The current estimate we have from our supplier for a time to repair is approximately 21h00 CEST (19h00 UTC). We will update this notice regularly as we obtain more information.
Our other circuits are unaffected.
12:30 UTC : Our supplier has indicated that they expect to be able to start initial repair splicing on the severed fibre within the next two hours.
14:45 UTC : Fibre engineers have begun installing replacement fibres in the affected section. They estimate approximately another three hours to complete resplicing work.
We are currently experiencing a problem with the toolbar display on a large number of websites created with Gandi Sitemaker. Out technichians are on location and working to fix the problem. We will keep you informed as to the status of the situation.
Please accept our apologies for any inconvienence this may cause.
UPDATE: This problem, due to an Adobe library update, has been fixed and the service is running normally again.
Due to a fiber cut in Paris, some domain registration activities may be unavailable or unresponsive as the registries concerned are not accessible via the internet or peering points. Connections via the PANAP and SFINX peering exchanges are especially affected. We are monitoring the situation and will update when we have more information. (Note that this fiber cut has an impact on many providers, and not only Gandi.)
Update: 10:10 CET: Connectivity restored for most peers at SFINX. Panap still has several peers down. Connectivity to registries should be operational now.
Due to a technical issue, DNS zone updates are not being processed since midday. Our technical teams are investigating the cause of the problem and are working to resolve the issue as soon as possible.
Update: 14:00 CET : The problem has been resolved.
We will be performing scheduled maintenance on all of our edge routers on Monday 16 August from 00:00 to 03:00 GMT (02:00 - 05:00 CET/Paris). During this window, each of our edge routers will be reloaded to upgrade the software with an expected outage of transit connectivity and peering on that router for up to fifteen minutes. After each router has been reloaded, we will wait for traffic to stabilise across all routers before performing the operation on the next router.
Customers should not encounter any significant impact to Gandi services other than a brief interruption of traffic on directly affected internet paths as the traffic is rerouted through other connections.
The Gandi Team
Updated: 06:30 16 Aug 2010: The maintenance has been completed and all routers are now operating normally.